OpenStudy (anonymous):

Find the x and y intercepts and graph the equation plooting the intercept......4x-7y=-28 show the work.. GET THE MEDAL

OpenStudy (aroub):

if u want to find the x-intercept u put y=0 and when u want to find the y-intercept u put x=0

OpenStudy (anonymous):

HUH?

OpenStudy (anonymous):

Let's try the first one. Let's substitute x=0 4(0) - 7y = -28 Solve for y and this will tell you the y intercept. (in other words where is y when x = 0)

OpenStudy (anonymous):

4x-7y+-28 4(0)-7y=-28 0-7y=-28 7y=-28 -- ---- 7 7 y=-4 can you tell me if i did it right

OpenStudy (anonymous):

Almost. Notice you had 0-7y=-28 but then the next line red 7y=-28 you need to be careful not to drop the negative :)

OpenStudy (anonymous):

where did I drop the negative

OpenStudy (anonymous):

When you went to divide both sides by 7. That should be dividing both sides by -7.

OpenStudy (anonymous):

o I see with the -7y

OpenStudy (anonymous):

so the answer will be a postive 4 in sead of a negative 4

OpenStudy (anonymous):

4x-7y=-28 4x-7(0)=-28 4x-0=-28 4x=-28 -- ---- 4 4 x=-7 right

OpenStudy (anonymous):

correct